# The connection of polaritonic chemistry with the physics of a spin glass

### v1

@@ Abstract
Polaritonic chemistry has garnered increasing attention in recent years due to pioneering experimental results, which show that site- and bond-selective chemistry at room temperature is achievable through strong collective coupling to field fluctuations in optical cavities. Despite these notable experimental strides, the underlying theoretical mechanisms remain unclear. In this focus review, we highlight a fundamental theoretical link between the seemingly unrelated fields of polaritonic chemistry and spin glasses, exploring its profound implications for the theoretical framework of polaritonic chemistry. Specifically, we present a mapping of the dressed electronic structure problem under collective vibrational strong coupling to the iconic Sherrington-Kirkpatrick model of spin glasses. This mapping uncovers a collectively induced instability in the dressed electronic structure (spontaneous replica symmetry breaking), which could provide the long-sought seed for significant local chemical modifications in polaritonic chemistry. This mapping paves the way to incorporate, adjust and probe numerous spin glass concepts in polaritonic chemistry, such as frustration, aging dynamics, excess of thermal fluctuations, time-reversal symmetry breaking or stochastic resonances. Ultimately, the mapping also offers fresh insights into the applicability of spin glass theory beyond condensed matter systems and it suggests novel theoretical directions such as polarization glasses with explicitly time-dependent order parameter functions.
@@ 1 Introduction
It is well established that molecular properties and chemical reactions can be influenced by light. Femtochemistry 1, 2 and coherent control using ultra-short and high-power lasers attest to this. So on a first glance it may seem straightforward to reach for a similar outcome by using optical cavities instead of laser driving, which has established the emergent field of polaritonic or QED chemistry (see Fig. 1). 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13 There are, however, a few very important differences that make polaritonic chemistry distinct. Firstly, when using lasers one tries to achieve site-selective chemistry usually with coherent, i.e., classical, light fields. In polaritonic chemistry usually a much smaller number of photons couples and their quantum nature becomes important. 3, 14 Secondly, in many cases the electromagnetic field inside an optical cavity is zero, i.e., the cavity is not pumped externally, such that only the strong coupling to the quantum and thermal fluctuations lead to modifications. 3, 14 Thirdly, and most importantly for this perspective, in most cases the coupling to a single molecule is small, but non-zero even in the thermodynamic limit. 3, 15, 16 Therefore, only the macroscopic ensemble of molecules a priori couples strongly to the photon-field fluctuations. This collective-coupling regime leads to a seemingly paradoxical situation: While chemical reactions and the properties of individual molecules are usually considered local in time and space, this traditional view is challenged in polaritonic chemistry due to strong feedback effects between the microscopic properties and the macroscopic behavior of the ensemble. The unique nature and origin of those feedback effects bridging different scales in time and space will the main focus of the present review.
@@
Connecting the different scales and isolating the physically relevant mechanism poses a formidable theoretical challenge, which has so far not been resolved satisfactorily due to its complexity (presumably its off-equilibrium glassy nature). 17, 5, 18, 19, 20 Particularly the origin of why in some molecular ensembles chemical reactions change 21, 22, 23, 24, 25, 26, 27 while in others under similar conditions no effect is observed,28, 29 remains elusive. In this article we want to provide a perspective on this conundrum that gives a partial answer and highlights a way forward to understand how the macroscopic behavior of an ensemble of molecules can act back on its individual constituents with the help of the electromagnetic modes of an optical cavity. While, as we will detail in this manuscript, there are many intricacies that need more theoretical and experimental investigations, eventually a relatively simple picture will emerge thanks to the established theoretical concepts of spin glasses. By borrowing ideas from this mature research discipline and applying them to the situation of collective vibrational strong coupling (VSC, i.e., the cavity is resonant to some vibrational degrees of freedom), we will see that the electronic polarizations of the individual molecules become synchronized via the cavity. However, since the overall macroscopic polarization of the ensemble is zero, these synchronized polarizations form intricate patterns that are (dynamically) frustrated. 30 That is, if the dynamic polarization of one molecule is flipped (or asynchronous), all the others need to react in a concerted synchronized way. In the following, we will call this situation a polarization glass20 The underlying electronic frustration effect does not only bridge length but also time scales. It acts back on all the other degrees of freedom such that it provides a necessary seed to trigger stochastic resonances, which together can affect reaction rates (rare events that depend on the details of the thermal ensemble and the local distribution of the energy in the ensemble) and other molecular properties. Indeed, based on the analogy with the spin glass, one expects non-trivial local and collective off-equilibrium effects even in a global thermal ensemble. Overall, the presented theoretical framework can provide an avenue for numerous future theoretical and experimental developments in the field of QED chemistry.
In the following we will first discuss the theoretical setting for describing VSC in the collective regime (Sec. 2), then show numerical and experimental evidence for frustration and phase-transition-like behavior of a polarization glass (Secs. 3.1-3.2), before we present a novel mapping of the dressed electronic-structure problem onto the prototypical Sherrington-Kirckpatrick (SK) spin glass model (Secs. 3.3-3.4) and introduce corresponding theoretical instability concepts (Secs. 3.5). We will also address theoretical and experimental observables as well as crucial differences between the spin-glass and polarization-glass setups (Sec. 4).
@@ 2 Pauli-Fierz ab initio theory
For simplicity we have chosen a single-effective cavity mode $\beta$, e.g., of a Fabry-Pérot cavity. Notice that more evolved cavity setups can be designed that, e.g., allow for higher mode-volume confinements (such as in plasmonic or micro cavities).32, 33, 34, 35, 36 Moreover, cavity leakage effects of the mirrors are a priori not captured by Eq. (1). However, those could be accounted for by considering multiple modes (broadening) and thus introducing a finite line width (lifetime) based on the imaginary part of the dielectric response of the mirrors.37, 38 The more general minimal-coupling Pauli-Fierz framework is discussed in, for example, Ref. 8, but we do not expect that this more intricate description will qualitatively change the results in the following.
@@
In Eq. (1) the free-space matter Hamiltonian is defined as $\hat{H}^{\rm m}$, which accounts for the quantized nuclei and electrons of the molecules. Thus $\hat{H}^{\rm m}$ describes the highly non-trivial free-space matter problem, which has been the focus of quantum-chemical methods over many decades. The second term describes the coupling of the matter to the quantized displacement field operator $\hat{q}_{\beta}$, with the conjugate photon operator defined as $\hat{p}_{\beta}$. The matter polarization operators for $N$ molecules with $N_{n}$ nuclei and $N_{e}$ electrons are given as $\hat{X}:=\boldsymbol{\lambda}\sum_{i=1}^{N}\sum_{n=1}^{N_{n}}Z_{n}\boldsymbol{\hat{R}}_{in}$ and $\hat{x}:=-\boldsymbol{\lambda}\sum_{i=1}^{N}\sum_{n=1}^{N_{e}}Z_{e}\boldsymbol{\hat{r}}_{in}$, where the nuclear and electronic total transition dipole moments, respectively, are coupled via $\boldsymbol{\lambda}$ to the effective displacement field mode of frequency $\omega_{\beta}$. The vectorial photon-matter coupling $\boldsymbol{\lambda}=\boldsymbol{\varepsilon}\lambda$ depends on the mode polarization vector $\boldsymbol{\varepsilon}$ and the coupling constant 8
@@
where $\mathscr{V}$ corresponds to the effective mode volume. This effective mode volume can be connected to properties of the Fabry-Pérot cavity and scales roughly as $L^{3}\mathscr{F}$, where $\mathscr{F}$ is the finesse of the cavity. A more elaborate discussion reveals that the effective mode volume leads to a finite light-matter coupling even in the macroscopic limit. 16. It is important to highlight two aspects of the Pauli-Fierz theory in the length gauge and the few-mode approximation. Firstly, that we have an equilibrium solution of the coupled system is due to the fact that the light-matter coupling term in Eq. (1) is quadratic and thus the Hamiltonian is bounded from below. Of specific importance for the stability of the coupled system is the term $(\hat{X}+\hat{x})^{2}$, which is quadratic in the coupling strength $\lambda$39, 40 This term is called dipole self-energy or self-polarization term in the literature. Secondly, for the proper free-space continuum limit it is important to subtract the free-space contributions from the effective-mode theory. Otherwise, one would double-count the interaction with the free-space modes that are captured by working with the observable masses of the charged particles. 16 Thus $\lambda=0$ means no cavity, and for any Fabry-Pérot cavity there is a finite mode volume $\mathscr{V}$ which implies $\lambda>0$ and also dictates the maximal amount of molecules that can in fact be coherently coupled for a given molecular density. This aspect will become important later when we discuss the scaling behavior in QED chemistry (see Sec. 3.5).
@@
Throughout this work, we are mainly interested in the physical properties of the cavity-mediated electronic structure given in Eq. (3). For this reason, we proceed by applying the classical (for the nuclear and displacement degrees of freedom) cavity Born-Oppenheimer approximation on the coupled nuclear-photon problem. This allows for a computationally efficient determination of reasonable parameters $(\boldsymbol{R},q_{\beta})$ that enter the dressed electronic-structure problem. In more detail, nuclei and (effective 16) displacement field evolve on the dressed ground-state electronic potential-energy surface according to the classical Hamiltonian dynamics of, 5, 20, 43, 44, 45, 46, 47, 48, 49
@@
The classical cavity Born-Oppenheimer approximation implies that any quantum and non-adiabatic effects of the nuclear structure are subsequently discarded, which allows for ground-state ab initio molecular dynamic implementations. Such a theoretical setup is numerically feasible even for large molecular ensembles $N\gg 1$. We will get back to this adiabatic assumption and discuss its validity in the context of VSC later in Secs. (3.5,4.1,4.2.3). Notice that assuming a classical displacement field $\boldsymbol{D}=\boldsymbol{\lambda}\omega_{\beta}q_{\beta}/4\pi$ does not mean that the transverse electric-field operator $\hat{\boldsymbol{E}}_{\perp}$ is entirely classical. Indeed, the transverse electric-field operator is given as, 20
@@
within the length-gauge representation used throughout this paper. Thus the electronic part of the macroscopic polarization operator $\hat{\boldsymbol{P}}=\boldsymbol{\lambda}(X+\hat{x})/4\pi$ remains fully quantized and describes the electronically bound photons of the hybrid light-matter states.39, 8 The quantized nature of $\hat{\boldsymbol{P}}$ will be an essential ingredient for all the subsequent discussions. A further important point is that the displacement field coordinate couples to the total dipole of the ensemble and the coupling scales linear in $\lambda$. Thus the coupling of the displacement field is different from the direct dipole-dipole coupling due to the self-energy term. Notice further that the longitudinal electric fields remain unaffected by the gauge choice, i.e., they correspond to the standard Coulomb interaction terms of the bare matter problem with classical nuclei and quantized electrons.

### v2

# Collectively-modified inter-molecular electron correlations: The connection of polaritonic chemistry and spin glass physics

@@ Abstract
Polaritonic chemistry has garnered increasing attention in recent years due to pioneering experimental results, which show that site- and bond-selective chemistry at room temperature is achievable through strong collective coupling to field fluctuations in optical cavities. Despite these notable experimental strides, the underlying theoretical mechanisms remain unclear. In this focus review, we highlight a fundamental theoretical link between the seemingly unrelated fields of polaritonic chemistry and spin glasses, exploring its profound implications for the theoretical framework of polaritonic chemistry. Specifically, we present a mapping of the dressed many-molecules electronic-structure problem under collective vibrational strong coupling to the spherical Sherrington-Kirkpatrick (SSK) model of a spin glass. This mapping uncovers a collectively induced instability of the intermolecular electron correlations, which could provide the long sought-after seed for significant local chemical modifications in polaritonic chemistry. Overall, the qualitative predictions made from the SSK model (e.g., dispersion effects, phase transitions, differently modified bulk and rare event properties, heating,…) agree well with available experimental observations. Our connection paves the way to incorporate, adjust and probe numerous spin glass concepts in polaritonic chemistry, such as modified fluctuation-dissipation relations, (non-equilibrium) aging dynamics, time-reversal symmetry breaking or stochastic resonances. Ultimately, the connection also offers fresh insights into the applicability of spin glass theory beyond condensed matter systems suggesting novel theoretical directions such as spin glasses with explicitly time-dependent (random) interactions.
@@ 1 Introduction
It is well established that molecular properties and chemical reactions can be influenced by light. Femtochemistry 1, 2 and coherent control using ultra-short and high-power lasers attest to this. So on a first glance it may seem straightforward to reach for a similar outcome by using optical cavities instead of laser driving, which has established the emergent field of polaritonic or QED chemistry (see Fig. 1). 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13 There are, however, a few very important differences that make polaritonic chemistry distinct and unique. Firstly, when using lasers one tries to achieve site-selective chemistry usually with coherent, i.e., classical, light fields. In polaritonic chemistry, usually a much smaller number of photons couples and their quantum nature becomes important. 3, 14 Secondly, in many cases the electromagnetic field inside an optical cavity is zero, i.e., the cavity is not pumped externally, such that only the strong coupling to the quantum and thermal fluctuations lead to modifications. 3, 14 Thirdly, and most importantly for this perspective, in most cases the coupling to a single molecule is small, but non-zero even in the thermodynamic limit. 3, 15, 16 Therefore, only the macroscopic ensemble of molecules a priori couples strongly to the photon-field fluctuations. This collective-coupling regime leads to a seemingly paradoxical situation: While chemical reactions and the properties of individual molecules are usually considered local in time and space, this traditional view is challenged in polaritonic chemistry due to strong feedback effects between the microscopic properties and the macroscopic behavior of the ensemble. The unique nature and origin of those feedback effects bridging different scales in time and space will the main focus of the present review.
@@
Connecting the different scales and isolating the physically relevant mechanism poses a formidable theoretical challenge, which has so far not been resolved satisfactorily due to its complexity (presumably its off-equilibrium glassy nature). 17, 5, 18, 19, 20 Particularly, the origin of why in some molecular ensembles chemical reactions change 21, 22, 23, 24, 25, 26, 27 while in others under similar conditions no effect is observed,28, 29 remains elusive. In this article we address the scaling conundrum by using established methods from spin glass physics. Those concepts provide a partial answer and highlight novel ways forward to understand how the macroscopic behavior of an ensemble of molecules can act back on its individual constituents with the help of the electromagnetic modes of an optical cavity. While, as we will detail in this manuscript, there are many intricacies that need more theoretical and experimental investigations, eventually a relatively simple picture will emerge thanks to the established theoretical concepts of spin glasses. By borrowing ideas from this mature research discipline and applying them to the situation of collective vibrational strong coupling (VSC, i.e., the cavity is resonant to some vibrational degrees of freedom), we will see that intermolecular electron correlations can possess spin glass features under certain conditions. An analogy to the spherical Sherrington-Kirkpatrick (SSK) model suggest not only a collective spin glass phase transition, but it also alters the local fluctuations (rare events) as well as the dynamics of the dressed electronic-structure problem. Resulting chemical consequences seem to agree with various experimental data. Moreover, the explicit time-dependence of the electron correlations (non-equilibrium aging dynamics) are expected to act back on all the other degrees of freedom such that it provides a seed to trigger chemically relevant stochastic resonance effects for the (ro)-vibrational degrees of freedom.30 Indeed, based on the analogy with the spin glass, one expects non-trivial local and collective off-equilibrium effects even in a global thermal ensemble. Overall, the presented theoretical framework can provide an avenue for numerous future theoretical and experimental developments in the field of QED chemistry.
In the following we will first discuss the theoretical setting for describing VSC in the collective regime and discuss how the cavity modifies statistical and thermodynamical considerations (Sec. 2). Then we show how intermolecular correlations lead to a connection to spin glass theory (Sec. 3). We then discuss some generic properties of spin glasses (Sec. 4), before highlighting potential consequences for polaritonic chemistry under collective VSC and comparing them to various experiments (Sec. 5). We briefly summarize our findings and future research perspectives in Sec. 6.
@@ 2 Pauli-Fierz ab initio theory
For simplicity we have chosen a single-effective cavity mode $\beta$, e.g., of a perfect (dissipationless) Fabry-Pérot cavity. Notice that more evolved cavity setups can be designed that, e.g., allow for higher mode-volume confinements (such as in plasmonic or micro cavities).32, 33, 34, 35, 36 Moreover, cavity leakage effects of the mirrors are a priori not captured by Eq. (1). However, those could be accounted for by considering multiple modes (broadening) and thus introducing a finite linewidth (lifetime) which leads in the continuum limit to the imaginary part of the dielectric response of the mirrors.37, 38 The more general minimal-coupling Pauli-Fierz framework is discussed in, for example, Ref. 8, but we do not expect that this more intricate description will qualitatively change the results in the following.
@@
In Eq. (1) the free-space matter Hamiltonian is defined as $\hat{H}^{\rm m}$, which accounts for the quantized nuclei and electrons of the molecules. Thus $\hat{H}^{\rm m}$ describes the highly non-trivial free-space matter problem, which has been the focus of quantum-chemical methods over many decades. The second term describes the coupling of the matter to the quantized displacement field operator $\hat{q}_{\beta}$, with the conjugate photon operator defined as $\hat{p}_{\beta}$. The matter polarization operators for $N$ molecules with $N_{n}$ nuclei and $N_{e}$ electrons are given as $\hat{X}:=\boldsymbol{\lambda}\cdot\sum_{i=1}^{N}\sum_{n=1}^{N_{n}}Z_{n}\boldsymbol{\hat{R}}_{in}$ and $\hat{x}:=-\boldsymbol{\lambda}\cdot\sum_{i=1}^{N}\sum_{n=1}^{N_{e}}Z_{e}\boldsymbol{\hat{r}}_{in}$, where the nuclear and electronic total transition dipole moments, respectively, are coupled via $\boldsymbol{\lambda}$ to the effective displacement field mode of frequency $\omega_{\beta}$. The vectorial photon-matter coupling $\boldsymbol{\lambda}=\boldsymbol{\varepsilon}\lambda$ depends on the mode polarization vector $\boldsymbol{\varepsilon}$ and the coupling constant 8
@@
where $\mathscr{V}$ corresponds to the effective mode volume. This effective mode volume can be connected to properties of the Fabry-Pérot cavity and scales roughly as $L^{3}\mathscr{F}$, where $\mathscr{F}$ is the finesse of the cavity. A more elaborate discussion reveals that the effective mode volume leads to a finite light-matter coupling even in the macroscopic limit. 16. It is important to highlight two aspects of the Pauli-Fierz theory in the length gauge and the few-mode approximation. Firstly, that we have an equilibrium solution of the coupled system is due to the fact that the light-matter coupling term in Eq. (1) is quadratic and thus the Hamiltonian is bounded from below. Of specific importance for the stability of the coupled system is the term $(\hat{X}+\hat{x})^{2}$, which is quadratic in the coupling strength $\lambda$39, 40 This term is called dipole self-energy or self-polarization term in the literature. Secondly, for the proper free-space continuum limit it is important to subtract the free-space contributions from the effective-mode theory. Otherwise, one would double-count the interaction with the free-space modes that are captured by working with the observable masses of the charged particles. 16 Thus $\lambda=0$ means no cavity, and for any Fabry-Pérot cavity there is a finite mode volume $\mathscr{V}$ which implies $\lambda>0$ and also dictates the maximal amount of molecules that can in fact be coherently coupled for a given molecular density. This aspect will become important later when we discuss the scaling behavior in QED chemistry (see Sec. 4.1.1).
@@
Throughout this work, we are mainly interested in the physical properties of the cavity-mediated electronic-structure given in Eq. (3). For this reason, we proceed by applying the classical (for the nuclear and displacement degrees of freedom) cavity Born-Oppenheimer approximation on the coupled nuclear-photon problem. This allows for a computationally efficient determination of reasonable parameters $(\boldsymbol{R},q_{\beta})$ that enter the dressed electronic-structure problem. In more detail, nuclei and (effective 16) displacement field evolve on the dressed ground-state electronic potential-energy surface according to the classical Hamiltonian dynamics of, 5, 20, 43, 44, 45, 46, 47, 48, 49
@@
The classical cavity Born-Oppenheimer approximation implies that any quantum and non-adiabatic effects of the nuclear structure are subsequently discarded, which allows for ground-state ab initio molecular dynamic implementations. Such a theoretical setup is numerically feasible even for large molecular ensembles $N\gg 1$. We will get back to this adiabatic assumption and discuss its validity in the context of VSC later in Secs. (4.1.1,4.1.2,4.1.4). Notice that assuming a classical displacement field $\boldsymbol{D}=\boldsymbol{\lambda}\omega_{\beta}q_{\beta}/4\pi$ does not mean that the transverse electric-field operator $\hat{\boldsymbol{E}}_{\perp}$ is entirely classical. Indeed, the transverse electric-field operator is given as, 20
@@
within the length-gauge representation used throughout this paper. Thus the electronic part of the macroscopic polarization operator $\hat{\boldsymbol{P}}=\boldsymbol{\lambda}(X+\hat{x})/4\pi$ remains fully quantized and describes the electronically bound photons of the hybrid light-matter states.39, 8 The quantized nature of $\hat{\boldsymbol{P}}$ will be an essential ingredient for all the subsequent discussions. A further important point is that the displacement field coordinate couples to the total dipole of the molecular ensemble and the coupling scales linear in $\lambda$. Thus the coupling of the displacement field is different from the direct dipole-dipole coupling due to the self-energy term. Notice further that the longitudinal electric fields remain unaffected by the gauge choice, i.e., they correspond to the standard Coulomb interaction terms of the bare matter problem with classical nuclei and quantized electrons.
2.1 Cavity Hartree-Fock (cHF) approximation
Assuming without loss of generality a system of $N$ identical molecules, each one possessing $N_{e}$ electrons, the cavity-Born-Oppenheimer Hartree-Fock (cBOHF) electronic wavefunction of the $(N\times N_{e})$-electron system is a Slater determinant of mutually orthonormal spin orbitals $\varphi_{i}$ 50:
$\Psi(\boldsymbol{\tau}_{11},\ldots\boldsymbol{\tau}_{N_{e}1},\ldots,\boldsymbol{\tau}_{N_{e}N})=\frac{1}{\sqrt{N_{e}!}}\langle\boldsymbol{\tau}_{11},\ldots,\boldsymbol{\tau}_{N_{e}1},\ldots\boldsymbol{\tau}_{N_{e}N}|\varphi_{11},\ldots,\varphi_{N_{e}1},\dots\varphi_{N_{e}N}\rangle\,.$ (7)
Here, $\boldsymbol{\tau}_{i}=(\boldsymbol{r}_{i}\sigma_{i})$ is used to denote the complete set of coordinates associated with the $i$-th electron, comprised of the spatial coordinate $\boldsymbol{r}_{i}$ and a spin coordinate $\sigma_{i}$. Note that the $(N\times N_{e})$-electron system described by $\Psi$ can be a single molecule (if $N=1$) or an ensemble of many molecules (if $N>1$). Thus, it is possible to treat cavity-induced interactions and standard Coulomb interactions in the same way. It is important to highlight that in the ensemble case we have besides intramolecular also intermolecular Coulomb and cavity-induced interactions. For the special case of the dilute gas limit, that is the electronic-structures of different molecules do not overlap, the ensemble Slater determinant may be replaced by a product 20 of individual molecular Slater determinants. Note that the displacement coordinate $q_{\beta}$ of the electric field mode is treated as a parameter in cBOHF ansatz, analogously to the nuclear coordinates, and thus is not part of the wave function. The resulting energy expression, where we have for the sake of brevity relabeled the electronic and nuclear coordinates as single sums, takes the form
$\bigl{\langle}\Psi\big{|}\hat{H}^{e}\big{|}\Psi\bigl{\rangle}$ $=$ $\bigl{\langle}\Psi\big{|}\sum_{i}^{N_{e}N}\bigg{\{}\frac{\hat{\boldsymbol{p}}_{i}^{2}}{2}-\frac{1}{2}\sum_{l}^{N_{n}N}\frac{Z_{l}}{|\hat{\boldsymbol{r}}_{i}-\boldsymbol{R}_{l}|}+\frac{1}{2}\sum_{j}^{N_{e}N}\frac{1}{|\hat{\boldsymbol{r}}_{i}-\hat{\boldsymbol{r}}_{j}|}\bigg{\}}+\bigg{(}\frac{1}{2}\hat{x}^{2}+\hat{x}X-\omega_{\beta}\hat{x}q_{\beta}\bigg{)}\big{|}\Psi\bigr{\rangle}$ (8)
[…]
$=$ $\underbrace{\big{\langle}\hat{h}^{\rm m}\big{\rangle}}_{=h^{\rm m}}+\underbrace{\big{\langle}\hat{h}^{\rm l}\big{\rangle}}_{=h^{\rm l}}+\underbrace{\big{\langle}\hat{J}_{\rm coul}\big{\rangle}}_{=J_{\rm coul}}+\underbrace{\big{\langle}\hat{K}_{\rm coul}\big{\rangle}}_{=K_{\rm coul}}+\underbrace{\big{\langle}\hat{J}_{\rm DSE}\big{\rangle}}_{=J_{\rm DSE}}+\underbrace{\big{\langle}\hat{K}_{\rm DSE}\big{\rangle}}_{=K_{\rm DSE}}$
where we used $\hat{h}^{\rm m}=\tfrac{\hat{\boldsymbol{p}}^{2}}{2}-\frac{1}{2}\sum_{l}^{N_{n}N}Z_{l}/(|\hat{\boldsymbol{r}}-\boldsymbol{R}_{l}|)$ and $\hat{h}^{\rm l}=\tfrac{1}{2}(Z_{e}\boldsymbol{\lambda}\cdot\hat{\boldsymbol{r}})^{2}-Z_{e}X\boldsymbol{\lambda}\cdot\hat{\boldsymbol{r}}+\omega_{\beta}q_{\beta}Z_{e}\boldsymbol{\lambda}\cdot\hat{\boldsymbol{r}}$. The two-electron integrals can be split into four different contributions: The Coulomb Hartree integral $J_{\rm coul}$, the Coulomb exchange integral $K_{\rm coul}$, the dipole-self energy (DSE) Hartree integral $J_{\rm DSE}$, and the DSE exchange integral $K_{\rm DSE}$. The standard variational procedure to search for the determinant $|\Psi\rangle$ with minimal energy $E_{0}=\langle\Psi_{0}|\hat{H}^{e}|\Psi_{0}\rangle$ is by variation of the spin orbitals $\phi_{i}$ such that $\delta\langle\Psi|\hat{H}^{e}|\Psi\rangle=0$, together with an orthogonality constraint $\langle\phi_{i}|\phi_{j}\rangle=\delta_{ij}$. This results in a non-linear eigenvalue problem for molecular orbitals that can be written in a compact form as
$\underbrace{\left[\hat{h}^{\rm m}+\hat{h}^{\rm l}+\hat{J}_{\rm coul}-\hat{K}_{\rm coul}+\hat{J}_{\rm DSE}-\hat{K}_{\rm DSE}\right]}_{=\hat{F}}\phi_{k}(\boldsymbol{\tau})=\epsilon_{k}\phi_{k}(\boldsymbol{\tau})$ (9)
Here the Fock operator $\hat{F}$ depends on the spin-orbitals $\phi_{k}(\boldsymbol{\tau})$ and is thus non-linear. This eigenvalue problem is usually computed iteratively.
Before we move on, we make an important observation about the accuracy of the Hartree-Fock method for a large ensemble of molecules or atoms. In the case of zero light-matter coupling, i.e., outside a cavity, it can be shown that the Hartree-Fock method becomes asymptotically exact as we approach the bulk/thermodynamic $N\rightarrow\infty$ limit 51, 52, 53.The reason for this is that the ground-state energy of the total ensemble scales as $N_{\rm tot}^{7/3}$, with $N_{\rm tot}=NN_{e}$ for charge neutral systems, while the difference between the exact and the Hartree-Fock solution scales as $N_{\rm tot}^{11/5}$. We therefore have that $\lim_{N_{\rm tot}\rightarrow\infty}(E_{\rm HF}[N_{\rm tot}]-E_{\rm exact}[N_{\rm tot}])/E_{\rm exact}[N_{\rm tot}]\rightarrow 0$. That is, the Hartree-Fock error compared to the total energy of the system vanishes. Such results are commonly used in the theory of phase transitions, where it is usually assumed that mean-field descriptions become asymptotically exact in the thermodynamic limit, i.e., for $N\rightarrow\infty$ 54. In this context it is also important to highlight that this argument does not imply that intramolecular properties of an individual molecule of the ensemble are treated exactly in this limit, but the behavior of the total ensemble is well represented by Hartree-Fock (or even Hartree) theory. This is why a statistical thermodynamic description of the ensemble of molecules becomes accurate 54. And when we consider chemical rates, they are as much a property of the individual constituent as of the total ensemble itself 55. Moreover, we need to highlight that for the behavior of an ensemble of molecules, the ground-state degeneracies and the amount of lowest-lying excited states are decisive 56. That is, the amount of available states (of the total ensemble) within an energy range set by the temperature, determines its phase and behavior. All of this will now be (at least slightly) modified by the cavity, which breaks the symmetry of free space and imposes a new time and length scale onto the molecular ensemble. Indeed, for the thermodynamic stability of the ensemble of molecules it is important to realize that the Coulomb-interaction gets screened and is short-ranged on an intermolecular scale 54, while the DSE terms are long-ranged due to their transverse nature and hence connect many more molecules. Let us elucidate these different points now in the following sections.
2.2 Dilute gas limit: cavity Hartree equations